OpenStudy (anonymous):

Which of the following is an equation that is parallel to the line y=3x+2, passing through the points (10,1) A.3x+y=29 B.3x-7=29 C. -3x-y=29 D.3x+y=29

OpenStudy (anonymous):

A line that is parallel will always have the same gradient as the other line, so your new line will start off like y=3x+c with m=3(same gradient) Then substitute the point given (10;1) to find C, as the graph passes through this point, So, 1=3(10) +c thus C=-29 so the equation is y=3x-29 just rearrange to get 3x-y =29 hope that helps :)